An object can possess kinetic energy, the energy of motion, and/or potential energy, the energy of position. Energy can be transferred in many ways, one of which is in the form of work. Work, w, is the energy transferred when an object is caused to move a distance, d, against a force, F. Mathematically, we would write this as

w=Fd

Force is the product of mass, m, and acceleration, a:

F=ma

When the acceleration is due to gravity, the symbol g is used instead of a, and its value is g = 9.81 m⋅s−2

How much work is done when a 125 g tomato is lifted 11.7 m ?

Express your answer with the appropriate units.
Answer #1

Solution -

Since the tomato is lifted so the work done to lift the tomato against the gravity has been stored in it in the form of potential energy so-

work done to lift the tomato = Increase in the potential energy of the tomatoes = m*g*h which is equal to -

125*10-3*9.81*11.7 = 14.34 Joules. (note - mass = 125*10-3 Kilograms )

So the work done to lift the tomatoes to height of 11.7 m is 14.34 Joules which is equal to increase in the potential energy of the tomatoes.

Please give a good rating if you finds the solution correct.

Thank you